OpenStudy (sidsiddhartha):

convolution of y(t)=exp(-3t)u(t)* u(t+3) is??

OpenStudy (sidsiddhartha):

having a problem to eliminate the step function from the integral can anyone help?

OpenStudy (accessdenied):

So you already put this into the form of an integral? If so, what did you have so far?

OpenStudy (sidsiddhartha):

L[exp(-3t)u(t)]=1/(s+3), L{u(t+3)}=exp(3s)/s so the out put is {1/(s+3)}*{exp(3s)/s}=exp(3s)/s(s+3) now i'm stuck what to do next?

OpenStudy (sidsiddhartha):

should i take inverse laplace now?@AccessDenied

OpenStudy (accessdenied):

Oh, yes. The convolution equals the inverse Laplace of the product of their Laplace transforms.

OpenStudy (sidsiddhartha):

hold on i'm doing it

OpenStudy (sidsiddhartha):

=1/3.exp(3s){1/s -1/(s+3)} =1/3{1-exp(-3t)}u(t+3) it is correct @AccessDenied

OpenStudy (sidsiddhartha):

i mean inverse laplace[1/3.exp(3s){1/s -1/(s+3)}]=1/3{1-exp(-3t)}u(t+3)

OpenStudy (sidsiddhartha):

is it okay @AccessDenied

OpenStudy (accessdenied):

Inverse Laplace of 1/3 exp(3s) (1/s - 1/(s+3)) = 1/3 inv.laplace(exp(3s) 1/s) - 1/3 inv.laplace(exp(3s) 1/(s+3)) Property: inverse laplace of e^(-cs) F(s) = u(t-c) f(t-c) So inverse laplace of exp(3s) 1/s) = u(t+3) *1 inverse laplace of exp(3s) 1/(s+3) = u(t+3) * e^(-3(t+3)) Is what I was getting

OpenStudy (sidsiddhartha):